OpenStudy (anonymous):

The sum of 3 consecutive odd integers is -21. Find the "middle" number. A. -9 B. -3 C. -7 D. -5

OpenStudy (anonymous):

-9, - 7, -5 sum to -21

OpenStudy (anonymous):

so B.?

OpenStudy (anonymous):

let first number=x second=x+1 third=x+2 x+x+1+x+2=-21 3x+3=-21 3x=-24 x=-8

OpenStudy (anonymous):

second= -8+1=-7 third=-8+2=-6

OpenStudy (anonymous):

@douglaswinslowcooper they are not consecutive -8 and -6 comes between them

OpenStudy (anonymous):

numbers are -8 -7 -6 middle is -7 @sadaholic19

OpenStudy (anonymous):

okay thanks for the correction @g17